OpenStudy (anonymous):

Help me plzz!! Medals Here!! Find the polynomial function with roots 1, -2, and 5.

OpenStudy (anonymous):

@amistre64 @ganeshie8 @iGreen @Jmmm @mathmale @ParthKohli @paki

OpenStudy (anonymous):

(x-1)(x+2)(x-5)=0 is the answer, however you may write in the form ax^3+bx^2+cx+d=0

OpenStudy (anonymous):

which is x^3-4x^2-7x+10=0

OpenStudy (anonymous):

could you help me with a couple more? @Princer_Jones Find the polynomial function with roots 1, 7, and -3 of multiplicity 2. and Find the polynomial function with roots 11 and 2i.

OpenStudy (anonymous):

The multiplicity of -3 is 2 or all of them has the multiplicity 2?

OpenStudy (anonymous):

All I believe

OpenStudy (anonymous):

let f(x)=(x-a)^k. Then the multiplicity of a is k. Use this formula and you yourself can compute easily.

OpenStudy (anonymous):

Is a,b,c are the roots of a polynomial, then (x-a)(x-b)(x-c) is the polynomial and i am sorry, in the previous question it was said to find the polynomial, so the 0 in the rhs will not be there in the answer

OpenStudy (anonymous):

*If